# Consistent eccentricities for gravitational wave astronomy: Resolving discrepancies between astrophysical simulations and waveform models

### v2

@@ III A consistent prescription for calculating eccentricity
where $\epsilon=\sqrt{1-e^{2}}$, and $\alpha_{i}$, $\beta_{i}$ are PN coefficients that depend on $e$ and $\eta$ Peters (1964); Junker and Schaefer (1992); Rieth and Schäfer (1997); Gopakumar and Iyer (1997); Arun et al. (2009) [^3] See Appendix A for explicit expressions of these.. For the purposes of this work, we will use equations accurate up to 2PN order (i.e. terms up to ($\order{v^{4}}$ )), and will neglect the effect of spin. We note that the full expressions are known up to 3PN order Arun et al. (2009) including the effect of spin Henry and Khalil (2023), but we find that restricting to 2PN gives results accurate enough for our purposes. Evolving the above equations in time starting from the initial conditions, we extract eccentricities at an appropriate reference frequency. This way of defining eccentricity makes it consistent with the one defined in waveforms, modulo systematics in modeling the GW emission itself. In the next section, we compare estimates derived from the prescription mentioned above (which we denote by $e_{t}^{\rm 2PN}$) to the those derived from the W03 prescription (denoted as $e_{\rm W03}$) as used in the astrophysical simulations literature, and also investigate potential biases due to incomplete physics.
@@
Figure 2: Comparison of orbital eccentricities in CMC simulations derived using the W03 prescription ($e_{\rm W03}$), to ones derived using the prescription in this work ($e_{t}^{\rm 2PN}$). The left panel shows a scatter plot of the two quantities. We see that there is considerable disagreement on the derived values of eccentricity, especially at $e\gtrsim 0.2$. This can be understood as the effect of the peak frequencies being shifted to higher $n_{\mathrm{peak}}$. The middle panel shows the histogram of (log) eccentricities for GW captures in globular clusters that occur during strong, few-body resonating interactions, derived using the W03 prescription and the prescription described in this work. The histograms do not deviate significantly from each other at low values of eccentricity, but are significantly different at $e\gtrsim 0.2$ (shaded region). The right panel shows the histogram of eccentricities in this $e\gtrsim 0.2$ region, where the discrepancies are more evident. Notably, the peak at $e\approx 1$ is no longer present in the histogram; this peak is an artifact of the procedure used to extract eccentricity in astrophysical simulations.
III.1 Comparison with the W03 prescription
@@
As an illustration of our prescription, we consider binary mergers simulated with the Cluster Monte Carlo (CMC) code Rodriguez et al. (2022). CMC performs long-term evolution of a globular cluster using a Hénon type Monte Carlo algorithm Hénon (1971a, b); Joshi et al. (2000, 2001); Fregeau et al. (2003); Fregeau and Rasio (2007); Chatterjee et al. (2010, 2013); Pattabiraman et al. (2013); Rodriguez et al. (2015) given a set of initial conditions for the cluster. While we will only consider the CMC simulations as an example in this section, we again note that the W03 prescription has been in many other works (see e.g. Samsing and Ramirez-Ruiz (2017); Samsing (2018); Rodriguez et al. (2018a); Zevin et al. (2019); Kritos et al. (2022); Chattopadhyay et al. (2023)).
@@
A catalog of CMC simulations on a grid of initial cluster masses, metallicity, radii, and galactocentric distances is publicly available Kremer et al. (2020, 2020). This catalog also includes information about compact object mergers that take place in these clusters. Merging compact binaries are extracted from these simulations using the following procedure:
@@
    During strong GW encounters, the approximated $N$-body dynamics (including leading-order PN corrections that account for orbital energy dissipation from GW emission Rodriguez et al. (2018b)) is evolved until two of the compact objects are in a bound state and reach a critical separation (typically $100M$), at which point the binary is assumed to definitively lead to a GW-driven merger and the properties of the binary are recorded.
@@
Instead of relying on the W03 definition of eccentricity, we apply the procedure outlined in Section III to each binary extracted from the simulations. We then calculate eccentricities at a fixed 22-mode reference frequency of 10 Hz. In the upper panel of Figure 2, we show a scatter plot of eccentricities derived from the two different definitions. Here, $e_{\rm W03}$ corresponds to the eccentricity as calculated by the W03 prescription, whereas $e_{t}^{\rm 2PN}$ corresponds to the one calculated using the prescription presented in this work. At low eccentricities, the two definitions agree with each other. However, as one increases the eccentricity, the disagreement between the two increases quite rapidly. This is a direct result of the peak frequency getting shifted to higher harmonics as one increases eccentricity; for higher eccentricity, a given peak frequency $f_{\rm peak,GW}$ will correspond to a lower $f_{22}$ (since $\frac{1}{2}n_{\rm peak}f_{22}=f_{\rm peak,GW}$), hence describing the $e_{t}^{\rm 2PN}$ at a much lower reference frequency. This also means that the binary will lose more of its eccentricity by the time it reaches $f_{\rm 22}=10\,{\rm Hz}$. As a consequence, $e_{t}^{\rm 2PN}$ is always less than $e_{\rm W03}$.
@@
Notably, all the points that have $e_{\rm W03}\approx 1$ have $e_{t}^{\rm 2PN}$ that range from $0.4-0.8$. We note that in many places in the literature, this $e\approx 1$ peak has been interpreted to be binaries that form “in-band” and merge quickly after getting captured; these are classically hyperbolic systems that are captured with a periapse frequency above the reference frequency considered (e.g., 10 Hz) via orbital energy loss through GW emission [^4] We note that the $e\approx 1$ value for systems in astrophysical simulations is somewhat artificial. These represent systems that do not have a defined eccentricity at a particular reference frequency according to the prescription outlined in Section II.2 and thus correspond to $e\geq 1$ at a particular reference frequency.. While that interpretation still remains qualitatively correct, our analysis shows that the value of eccentricity that should be supplied to waveforms in comparison studies is not $e\approx 1$ but rather is shifted to much lower values.
@@
The middle panel of Figure 2 shows histograms of $e_{\rm W03}$ and $e_{t}^{\rm 2PN}$ calculated from the CMC simulations, with appropriate weights applied to each sample based on cluster mass, metallicity, and detection probability [^5] See e.g. Appendix A of Ref. Zevin et al. (2021b) (and references therein) for a description of this weighting procedure.. As would be expected from the preceding discussions, the histograms match for low values of eccentricity, and start deviating at $e\gtrsim 0.2$. Most strikingly, the peak at $e_{\rm W03}\approx 1$ in the $e_{\rm W03}$ histogram is not present in the $e_{t}^{\rm 2PN}$ histogram. These observations are more evident in the right panel of Figure 2, where we show the histograms of eccentricities for $e>0.2$. Since the waveform changes significantly as a function of eccentricity, one would expect these discrepancies to impact any analysis that assumes consistency between the definitions of eccentricities in waveforms and astrophysical simulations. We will comment on this briefly in Section III.4.
@@
It is natural to ask whether the discrepancy is primarily driven by the mismatch in conventions for the reference frequency, or due to only accounting for the leading order PN term in the W03 prescription. The results in the preceding paragraphs would lead us to believe that it is the former. We test this in another way: since we know $n_{\rm peak}(e)$, we can also adjust the eccentricities calculated via W03 as a post-processing step by assuming that the $e_{\rm W03}$ is defined at a reference frequency $f_{\rm 22}=2\times f_{\rm peak}/n_{\rm peak}(e)$, and evolve to a fixed reference (22-mode) frequency using PN equations (Eqs. (8)-(10)) but only up to the leading order. Let the eccentricity calculated using this modified version of the W03 procedure be $e_{\rm W03}^{f_{22}}$. We find that $e_{\rm W03}^{f_{22}}\approx e_{t}^{\rm 2PN}$ for the binaries we consider [^6] We exclude binaries with $e_{\rm W03}\approx 1$ from this calculation, due to their aforementioned artificial nature., with a $\sim 5\%$ scatter that could be attributed to non-inclusion of higher-order PN terms (see also Section III.2). This further confirms that the discrepancy in the calculated eccentricities is primarily driven by the difference in conventions for the reference frequency.
Figure 4: The top panel compares $e_{t}^{\rm 0PN}$ and $e_{t}^{\rm 2PN}$ computed at $f_{\mathrm{ref}}=20\,\rm Hz$ for three initial frequencies $f_{0}=5,10,15\,{\rm Hz}$ as a function of initial eccentricity $e_{0}$. The bottom panel makes a similar comparison between $e_{t}^{\rm 2PN}$ calculated assuming a value $\chi_{z}=\chi_{1,z}=\chi_{2,z}$, and $\chi_{z}=0$. We choose two representative values of $\chi_{z}$: $\chi_{z}=0.7$ (dotted lines), and $\chi_{z}=0.99$ (dashed lines).
III.2 Effects of ignoring higher order PN terms and spin effects
We now systematically investigate if there is a relative advantage in using the 2PN equations above just the leading order (0PN) evolution equations. For an equal mass binary of total mass $M=50\,M_{\odot}$, we consider three different starting frequencies $f_{0}=5,10,15\,{\rm Hz}$, and calculate $e_{t}^{\rm 2PN}$ and $e_{t}^{\rm 0PN}$ at a reference frequency of 20 Hz for a range of initial eccentricities $e_{0}$. The fractional differences between $e_{t}^{\rm 2PN}$ and $e_{t}^{\rm 0PN}$ estimates are plotted in the top panel of Figure 4, and show an increasing trend for lower values of $f_{0}$, with a discrepancy as high as $12\%$ for $f_{0}=5\,{\rm Hz}$. For a given $f_{0}$, the fractional differences are lower for higher $e_{0}$. In typical applications, one might want to calculate eccentricities at a given reference frequency for much lower values of $f_{0}$. While a $\sim 10\%$ difference might not be relevant while comparing eccentricity estimates from data to simulations, we recommend using $e_{t}^{\rm 2PN}$ estimates since the costs of evaluating $e_{t}^{\rm 0PN}$ and $e_{t}^{\rm 2PN}$ are similar (see Appendix C).
@@ III.2 Effects of ignoring higher order PN terms and spin effects
III.3 Comparison to estimates derived from gw_eccentricity package
Figure 5: Comparison between the eccentricity definition given by Ref. Shaikh et al. (2023), calculated using the waveform approximant SEOBNRv4EHM ($e_{\rm gw}^{\rm SEOB}$; dotted lines), and $e_{t}^{\rm 2PN}$ (dashed lines), as a function of the initial eccentricity, $e_{0}$, for three initial frequencies, $f_{0}=5,10,15\,{\rm Hz}$. We assume an equal mass non-spinning binary with total mass $M=50\,M_{\odot}$ for this purpose.
As mentioned earlier, different GW waveform approximants have different conventions to define eccentricity. To solve this issue, Ref. Shaikh et al. (2023) proposed a standardized definition of eccentricity that can be applied to any waveform approximant, $e_{\rm gw}$, which is given by
@@ III.3 Comparison to estimates derived from gw_eccentricity package
and $\omega_{22}^{\rm p,a}(t)$ are the 22-mode frequencies through the pericenters and apocenters, respectively. This standardized definition is proposed in order to extract eccentricity directly from observables, in this case the gravitational waveform. More details on the motivations for this definition and the various methods of calculating $\omega_{22}^{\rm p,a}(t)$ can be found in Ref. Shaikh et al. (2023). We also briefly investigate the differences between $e_{\rm gw}$ derived from different waveform approximants in Appendix B.
@@
While $e_{\rm gw}$ does asymptote to the PN definition of eccentricity, these two definitions may differ in general. For instance, waveforms could include more accurate dynamics at high eccentricity and closer to merger as compared to our 2PN prescription. Instead of the 2PN accurate equations, we could hence imagine replacing the prescription in Section III by generating waveforms at the specified initial epochs, and let gw_eccentricity calculate the eccentricities at the required reference frequency.
@@
However, since calculating $e_{\rm gw}$ involves computing the waveform starting at a specified initial frequency/time, the calculation would be on the order of a hundred times slower (see Appendix C) than computing the PN eccentricity. The cost of generating waveforms also increases with decreasing starting frequency. Since astrophysical simulations require calculating eccentricites of a large number of binaries, calculating the Ref. Shaikh et al. (2023) definition of eccentricity would be computationally inefficient. Therefore, we now compare the differences between the two eccentricity definitions, $e_{\rm gw}$ and $e_{t}^{\rm 2PN}$, and quantify the reliability of $e_{t}^{\rm 2PN}$ eccentricity estimates from astrophysical simulations. For the investigations that follow, we will use the SEOBNRv4EHM Ramos-Buades et al. (2022) waveform approximant for the calculation of $e_{\rm gw}$. Using the same mass, $f_{0}$, and $e_{0}$ values as in Section III.2, we calculate $e_{t}^{\rm 2PN}$ and $e_{\rm gw}^{\rm SEOB}$ at a reference frequency of 20 Hz. Figure 5 illustrates the differences between these. While small (a few tens of a percent at worst), the differences do accumulate as the binary evolves for a longer period of time with $f_{0}$ values closer to the reference frequency giving more agreement. Since the cost of computing $e_{\rm gw}$ is a factor of $\sim 100$ greater as compared to calculating $e_{t}^{\rm 2PN}$ (see Appendix C), we conclude that $e_{t}^{\rm 2PN}$ is sufficient, especially in regimes where PN theory is valid.
III.4 Implications for existing work

### v3

@@ 3 A consistent prescription for calculating eccentricity
where $\epsilon=\sqrt{1-e^{2}}$, and $\alpha_{i}$, $\beta_{i}$ are PN coefficients that depend on $e$ and $\eta$ (Peters, 1964; Junker & Schaefer, 1992; Rieth & Schäfer, 1997; Gopakumar & Iyer, 1997; Arun et al., 2009) [^3] See Appendix A for explicit expressions of these.. For the purposes of this work, we will use equations accurate up to 2PN order (i.e. terms up to ($\order{v^{4}}$ )), and will neglect the effect of spin. We note that the full expressions are known up to 3PN order (Arun et al., 2009) including the effect of spin (Henry & Khalil, 2023), but we find that restricting to 2PN gives results accurate enough for our purposes [^4] We could also equivalently use the approximate analytical expression generalizing Eq. (7) derived by Tucker & Will (2021) instead of evolving the PN differential equations, but we concentrate on the latter in this work.. Evolving the above equations in time starting from the initial conditions, we extract eccentricities at an appropriate reference frequency. This way of defining eccentricity makes it consistent with the one defined in waveforms, modulo systematics in modeling the GW emission itself. In the next subsection, we compare estimates derived from the prescription mentioned above (which we denote by $e_{t}^{\rm 2PN}$) to those derived from the W03 prescription (denoted as $e_{\rm W03}$) as used in the astrophysical simulations literature, and also investigate potential biases due to incomplete physics.
@@
Figure 2: Comparison of orbital eccentricities in CMC simulations derived using the W03 prescription ($e_{\rm W03}$), to ones derived using the prescription in this work ($e_{t}^{\rm 2PN}$) at a reference frequency of $10\,{\rm Hz}$ set by each definition. The left panel shows a scatter plot of the two quantities. We see that there is considerable disagreement on the derived values of eccentricity, especially at $e\gtrsim 0.2$. This can be understood as the effect of the peak frequencies being shifted to higher $n_{\mathrm{peak}}$. The middle panel shows the histogram of (log) eccentricities for GW captures in globular clusters that occur during strong, few-body resonating interactions, derived using the W03 prescription and the prescription described in this work. The histograms do not deviate significantly from each other at low values of eccentricity, but are significantly different at $e\gtrsim 0.2$ (shaded region). The right panel shows the histogram of eccentricities in this $e\gtrsim 0.2$ region, where the discrepancies are more evident. Notably, the peak at $e\approx 1$ is no longer present in the histogram; this peak is an artifact of the procedure used to extract eccentricity in astrophysical simulations.
3.1 Comparison with the W03 prescription
@@
As an illustration of our prescription, we consider binary mergers simulated with the Cluster Monte Carlo (CMC) code (Rodriguez et al., 2022). CMC performs long-term evolution of a globular cluster using a Hénon type Monte Carlo algorithm (Hénon, 1971a, b; Joshi et al., 2000, 2001; Fregeau et al., 2003; Fregeau & Rasio, 2007; Chatterjee et al., 2010, 2013; Pattabiraman et al., 2013; Rodriguez et al., 2015) given a set of initial conditions for the cluster. While we will only consider the CMC simulations as an example in this section, we again note that the W03 prescription has been in many other works (see e.g. Samsing & Ramirez-Ruiz 2017; Samsing 2018; Rodriguez et al. 2018a; Zevin et al. 2019; Kritos et al. 2022; Chattopadhyay et al. 2023).
@@
A catalog of CMC simulations on a grid of initial cluster masses, metallicity, radii, and galactocentric distances is publicly available (Kremer et al., 2020, 2020). This catalog also includes information about compact object mergers that take place in these clusters. Merging compact binaries are extracted from these simulations using the following procedure:
@@
    During strong GW encounters, the approximated $N$-body dynamics (including leading-order PN corrections that account for orbital energy dissipation from GW emission (Rodriguez et al., 2018b)) is evolved until two of the compact objects are in a bound state and reach a critical separation (typically $100M$), at which point the binary is assumed to definitively lead to a GW-driven merger and the properties of the binary are recorded.
@@
Instead of relying on the W03 definition of eccentricity, we apply the procedure outlined in Section 3 to each binary extracted from the simulations. We then calculate eccentricities at a fixed 22-mode reference frequency of 10 Hz. In the left panel of Figure 2, we show a scatter plot of eccentricities derived from the two different definitions. Here, $e_{\rm W03}$ corresponds to the eccentricity as calculated by the W03 prescription, whereas $e_{t}^{\rm 2PN}$ corresponds to the one calculated using the prescription presented in this work. At low eccentricities, the two definitions agree with each other. However, as one increases the eccentricity, the disagreement between the two increases quite rapidly. This is a direct result of the peak frequency getting shifted to higher harmonics as one increases eccentricity; for higher eccentricity, a given peak frequency $f_{\rm peak,GW}$ will correspond to a lower $f_{22}$ (since $\frac{1}{2}n_{\rm peak}f_{22}=f_{\rm peak,GW}$), hence describing the $e_{t}^{\rm 2PN}$ at a much lower reference frequency. This also means that the binary will lose more of its eccentricity by the time it reaches $f_{\rm 22}=10\,{\rm Hz}$. As a consequence, $e_{t}^{\rm 2PN}$ is always less than $e_{\rm W03}$.
@@
Notably, all the points that have $e_{\rm W03}\approx 1$ have $e_{t}^{\rm 2PN}$ that range from $0.4-0.8$. We note that in many places in the literature, this $e\approx 1$ peak has been interpreted to be binaries that form “in-band” and merge quickly after getting captured; these are classically hyperbolic systems that are captured with a periapse frequency above the reference frequency considered (e.g., 10 Hz) via orbital energy loss through GW emission [^5] We note that the $e\approx 1$ value for systems in astrophysical simulations is somewhat artificial. These represent systems that do not have a defined eccentricity at a particular reference frequency according to the prescription outlined in Section 2.2 and thus correspond to $e\geq 1$ at a particular reference frequency.. While that interpretation still remains qualitatively correct, our analysis shows that the value of eccentricity that should be supplied to waveforms in comparison studies is not $e\approx 1$ but rather is shifted to much lower values.
@@
The middle panel of Figure 2 shows histograms of $e_{\rm W03}$ and $e_{t}^{\rm 2PN}$ calculated from the CMC simulations, with appropriate weights applied to each sample based on cluster mass, metallicity, and detection probability [^6] See e.g. Appendix A of Zevin et al. (2021a) (and references therein) for a description of this weighting procedure.. As would be expected from the preceding discussions, the histograms match for low values of eccentricity, and start deviating at $e\gtrsim 0.2$. Most strikingly, the peak at $e_{\rm W03}\approx 1$ in the $e_{\rm W03}$ histogram is not present in the $e_{t}^{\rm 2PN}$ histogram. These observations are more evident in the right panel of Figure 2, where we show the histograms of eccentricities for $e>0.2$. Since the waveform changes significantly as a function of eccentricity, one would expect these discrepancies to impact any analysis that assumes consistency between the definitions of eccentricities in waveforms and astrophysical simulations. We will comment on this briefly in Section 3.4.
@@
It is natural to ask whether the discrepancy is primarily driven by the mismatch in conventions for the reference frequency, or due to only accounting for the leading order PN term in the W03 prescription. The results in the preceding paragraphs would lead us to believe that it is the former. We test this in another way: since we know $n_{\rm peak}(e)$, we can also adjust the eccentricities calculated via W03 as a post-processing step by assuming that the $e_{\rm W03}$ is defined at a reference frequency $f_{\rm 22}=2\times f_{\rm peak}/n_{\rm peak}(e)$, and evolve to a fixed reference (22-mode) frequency using PN equations (Eqs. (8)-(10)) but only up to the leading order. Let the eccentricity calculated using this modified version of the W03 procedure be $e_{\rm W03}^{f_{22}}$. We find that $e_{\rm W03}^{f_{22}}\approx e_{t}^{\rm 2PN}$ for the binaries we consider [^7] We exclude binaries with $e_{\rm W03}\approx 1$ from this calculation, due to their aforementioned artificial nature., with a $\sim 5\%$ scatter that could be attributed to non-inclusion of higher-order PN terms (see also Section 3.2). This further confirms that the discrepancy in the calculated eccentricities is primarily driven by the difference in conventions for the reference frequency.
Figure 4: The top panel compares $e_{t}^{\rm 0PN}$ and $e_{t}^{\rm 2PN}$ computed at $f_{\mathrm{ref}}=20\,\rm Hz$ for three 22-mode starting frequencies $f_{0}=5,10,15\,{\rm Hz}$ as a function of initial eccentricity $e_{0}$. The bottom panel makes a similar comparison between $e_{t}^{\rm 2PN}$ calculated assuming a value $\chi_{z}=\chi_{1,z}=\chi_{2,z}$, and $\chi_{z}=0$. We choose two representative values of $\chi_{z}$: $\chi_{z}=0.7$ (dotted lines), and $\chi_{z}=0.99$ (dashed lines).
3.2 Effects of ignoring higher order PN terms and spin effects
We now systematically investigate if there is a relative advantage in using the 2PN equations above just the leading order (0PN) evolution equations. For an equal mass binary of total mass $M=50\,M_{\odot}$, we consider three different 22-mode starting frequencies $f_{0}=5,10,15\,{\rm Hz}$, and calculate $e_{t}^{\rm 2PN}$ and $e_{t}^{\rm 0PN}$ at a reference frequency of 20 Hz for a range of initial eccentricities $e_{0}$. The fractional differences between $e_{t}^{\rm 2PN}$ and $e_{t}^{\rm 0PN}$ estimates are plotted in the top panel of Figure 4, and show an increasing trend for lower values of $f_{0}$, with a discrepancy as high as $12\%$ for $f_{0}=5\,{\rm Hz}$. For a given $f_{0}$, the fractional differences are lower for higher $e_{0}$. In typical applications, one might want to calculate eccentricities at a given reference frequency for much lower values of $f_{0}$. While a $\sim 10\%$ difference might not be relevant while comparing eccentricity estimates from data to simulations, we recommend using $e_{t}^{\rm 2PN}$ estimates since the costs of evaluating $e_{t}^{\rm 0PN}$ and $e_{t}^{\rm 2PN}$ are similar (see Appendix C).
@@ 3.2 Effects of ignoring higher order PN terms and spin effects
3.3 Comparison to estimates derived from gw_eccentricity package
Figure 5: Comparison between the eccentricity definition given by Shaikh et al. (2023), calculated using the waveform approximant SEOBNRv4EHM ($e_{\rm gw}^{\rm SEOB}$; dotted lines), and $e_{t}^{\rm 2PN}$ (dashed lines), as a function of the initial eccentricity, $e_{0}$, for three initial frequencies, $f_{0}=5,10,15\,{\rm Hz}$. We assume an equal mass non-spinning binary with total mass $M=50\,M_{\odot}$ for this purpose.
As mentioned earlier, different GW waveform approximants have different conventions to define eccentricity. To solve this issue, Shaikh et al. (2023) proposed a standardized definition of eccentricity that can be applied to any waveform approximant, $e_{\rm gw}$, which is given by
@@ 3.3 Comparison to estimates derived from gw_eccentricity package
and $\omega_{22}^{\rm p,a}(t)$ are the 22-mode frequencies through the pericenters and apocenters, respectively. This standardized definition is proposed in order to extract eccentricity directly from observables, in this case the gravitational waveform. More details on the motivations for this definition and the various methods of calculating $\omega_{22}^{\rm p,a}(t)$ can be found in Shaikh et al. (2023). We also briefly investigate the differences between $e_{\rm gw}$ derived from different waveform approximants in Appendix B.
@@
While $e_{\rm gw}$ does asymptote to the PN definition of eccentricity in the small eccentricity limit as well as the large separation limit, these two definitions may differ in general. For instance, waveforms could include more accurate dynamics at high eccentricity and closer to merger as compared to our 2PN prescription. Instead of the 2PN accurate equations, we could hence imagine replacing the prescription in Section 3 by generating waveforms at the specified initial epochs, and let gw_eccentricity calculate the eccentricities at the required reference frequency.
@@
However, since calculating $e_{\rm gw}$ involves computing the waveform starting at a specified initial frequency/time, the calculation would be on the order of a hundred times slower (see Appendix C) than computing the PN eccentricity. The cost of generating waveforms also increases with decreasing starting frequency. Since astrophysical simulations require calculating eccentricites of a large number of binaries, calculating the Shaikh et al. (2023) definition of eccentricity would be computationally inefficient. Therefore, we now compare the differences between the two eccentricity definitions, $e_{\rm gw}$ and $e_{t}^{\rm 2PN}$, and quantify the reliability of $e_{t}^{\rm 2PN}$ eccentricity estimates from astrophysical simulations. For the investigations that follow, we will use the SEOBNRv4EHM (Ramos-Buades et al., 2022) waveform approximant for the calculation of $e_{\rm gw}$. Using the same mass, $f_{0}$, and $e_{0}$ values as in Section 3.2, we calculate $e_{t}^{\rm 2PN}$ and $e_{\rm gw}^{\rm SEOB}$ at a reference frequency of 20 Hz. Figure 5 illustrates the differences between these. While small (a few tens of a percent at worst), the differences do accumulate as the binary evolves for a longer period of time with $f_{0}$ values closer to the reference frequency giving more agreement. Since the cost of computing $e_{\rm gw}$ is a factor of $\sim 100$ greater as compared to calculating $e_{t}^{\rm 2PN}$ (see Appendix C), we conclude that $e_{t}^{\rm 2PN}$ is sufficient, especially in regimes where PN theory is valid.
3.4 Implications for existing work